OpenStudy (anonymous):

The basketball team scored 12 points the first game and 47 points the eighth game of the season. How many point will they earn with this pattern on the fourteenth game of the season? answers to choose from. a)68 b)71 c)74 d)77

OpenStudy (anonymous):

@T-Wrx

OpenStudy (anonymous):

@bloodrain

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

Can you help me

OpenStudy (anonymous):

Looks like it is in A.P with first term 12 and common difference 5

OpenStudy (anonymous):

So, can u find t14?

OpenStudy (anonymous):

ive already tried but cant seem to get it

OpenStudy (anonymous):

t14 = a1 + (14-1)d =12 + 13*5 =77

OpenStudy (anonymous):

thank you soo much so much

OpenStudy (anonymous):

welcome